PROBLEM1 You have been placed in charge of campground development for your agency. You have been asked to assess the value of two potential options using a discount rate is 3% (annual). 1a) Upgraded facilities (hookups and restrooms) will require an initial investment of $65,000 now and require maintenance costs of $1500 per year starting at the end of the first year and $23,000 every five years starting at the end of the fifth year. You expect 750 overnight visitors each year and will charge $35 per night. What is the Present Value of this plan in perpetuity? 1b) Rustic facilities (no hookups and outhouses) will require an initial investment of $15,000 now and require maintenance costs of $100 per year starting at the end of the first year an additional $3,000 every three years starting at the end of the third year. You expect 1750 overnight visitors each year and will charge $12 per night. What is the Present Value of this plan in perpetuity? 1c) Which option would you prefer?
Answer #1

Answer:

1a)

Given

Initial Investment P=$65000

Maintenance cost M=$1300 per year

Maintenance cost N=$23000 every five year

Overnight visitors each year= 750

Price per night=$35

So total revenue TR=35*750=$26250

discount rate r=13%

PV of maintenance cost for 5 years PVM=M*(1-(1+r)^-5)/r+ N/(1+r)^5=1300*(1-(1+13%)^-5)/13% + 23000/(1+13%)^5=$17055.88

So NPV of this plan = TR/r - PVM*(1+r)/r-P=26250/13%-17055.88*(1+13%)/13% -65000=-$11331.83 Eq1

1b)

Given

Initial Investment P=$15000

Maintenance cost M=$100 per year

Maintenance cost N=$3000 every three year

Overnight visitors each year= 1750

Price per night=$12

So total revenue TR=12*1750=$21000

discount rate r=13%

PV of maintenance cost for 5 years PVM=M*(1-(1+r)^-3)/r+ N/(1+r)^3=100*(1-(1+13%)^-3)/13% + 3000/(1+13%)^3=$1980

So NPV of this plan = TR/r - PVM*(1+r)/r-P=21000/13%-1980*(1+13%)/13% -15000=$129327.68 Eq 3

1C) from equation 1 and 2 we find the PV for no hookups and outhouse is greater than plan with  hookups and outhouse. So we will prefer plan with no hookups and outhouse.